OpenStudy (anonymous):

Select the equation of the line that passes through the point (3, 5) and is perpendicular to the line x = 4. y = 5 x = 5 y = 4 x = 3

OpenStudy (anonymous):

@misssunshinexxoxo

OpenStudy (anonymous):

@keigh2015

OpenStudy (misssunshinexxoxo):

Do you know point slope form?

OpenStudy (anonymous):

y-y=m(x-x)

OpenStudy (misssunshinexxoxo):

Plug it in

OpenStudy (anonymous):

okay so y-5=m(x-3)?

OpenStudy (misssunshinexxoxo):

Yes

OpenStudy (keigh2015):

I think it is y=4 I already did this test but I can't get to it because I finished the class already.

OpenStudy (anonymous):

ooh, what does the word perpendicular mean again?

OpenStudy (misssunshinexxoxo):

Hello Keigh kindly please refrain from posting direct answers; rather please help the person understand

OpenStudy (misssunshinexxoxo):

@paki

OpenStudy (keigh2015):

yeah i am going to help her with it once i remember how to do it exactly

OpenStudy (keigh2015):

Perpendicular means when two lines cross over each other like take your arms and make a plus sign with them

OpenStudy (misssunshinexxoxo):

\[Y =m(x-3) + 5\]

OpenStudy (anonymous):

okay thanks

OpenStudy (anonymous):

so now would i solve that? or

OpenStudy (misssunshinexxoxo):

Yes

OpenStudy (anonymous):

y=4 ?

OpenStudy (misssunshinexxoxo):

Yes way to go! :)

OpenStudy (anonymous):

thank you so much !

OpenStudy (anonymous):

can i ask one quick question

OpenStudy (anonymous):

esto es español no matemáticas

OpenStudy (anonymous):

@DezeliA yo se pero lo puse en de mal puesto desculpa